1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
Varulven  Le loup-garou  The Werewolf  Der Werwolf
tradukita de Sven Collberg  tradukita de R. Platteau  tradukita de Alexander Gross

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konju-

gera.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle !

”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich be-

gab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge

mich!“

mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på

pass med tåligt korslagd tass på

tass.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle !

”

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then

was ready.

Der Dorfschulmeister stieg hin-

auf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem Toten:

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv.

Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“ - On peut à ton sujet puisque nous y som-

mes Se demander encor bien des choses en som-

me. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sen-

se.”

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“

…  …  …  …

2/3

Varulven myste gladåt hur pers. hade följts av sing. och

plur. “Fast”, bad han, “kan det inte

ges ens en enda pers. i tempus pre-

sens?”

Loup-garcomment ? et puis

Bien évidemment loup-garqui ? ”

Toutes ces questions semblèrent au loup bien

absconses, Car en fait ce qu’il voulait c’était les

réponses.

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“

Bekänna måste vår magister, att därvidlag fanns vissa bri-

ster. En varulv sine imperfecto var contradictio in adjecto.

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moitié “A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utter-

ly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinular.

Varulven satt med tårad lins. Han visste att han fanns och

finns. Därtill kom också fru m.m. som även måste existera.

Ne pourrons plus me faire me sentir tout en-

tier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

The Werewolf knew better–his sons still

slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Der Wolf erhob sich tränenblind

– er hatte ja doch Weib und Kind! Doch da er keine Gelehrter

eben, so schied er dankend und erge-

ben.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow.

com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)